The first term of sequnce is x. To get the next term multiply the previous term by 2 and add 1.the third term of the sequence is 21.find the value of x. Plzz help me find the answer...its 4.5 in the book gcse
Answer by JulietG(1812)

Your sequence is:
(x, 2x + 1, 21)
You can work this backward.
Find the middle number first.
21 - 1 = 2x
20 = 2x
10 = x
.
Now take the 10 and work backward with the same formula.
10 - 1 = 2x
9 = 2x
4.5 = x
.
Let's plug it in to make certain.
4.5 * 2 = 9, plus 1 = 10
10 * 2 = 20, plus 1 = 21
Success!